Lassie, a very energetic dog with lots of stamina, is walking along a country road. She is walking 3 miles per hour. 6 hours into her journey, Lassie is at mile marker 100. What mile marker will Lassie be at 7 hours into her journey? How about 106 hours into her journey? (Assume Lassie is an extraordinary dog and never needs to stop for a rest.) What mile marker will Lassie be “n” hours into her journey?
Answer by richwmiller(17219)

6,100
100 = 82 + (6*3)
7 hours 103
a7 = 82 + 3*7
a7=103
an=82+3n